# Quantum Edge Detection

### v1

@@ 1 Introduction
In a recent paper [19], the concept of unsupervised classification of quantum data was introduced, focusing on systems with unknown quantum states. The paper presents an optimal single-shot protocol for binary classification, capable of automatically sorting disordered input arrays with minimal assumptions while preserving partial information. It is demonstrated that this protocol surpasses classical methods for dimensionality three or higher. Furthermore, the quantum protocol provides efficient execution on quantum computers, in stark contrast to the NP-hard complexity of classical approaches. Its potential applications in high-energy physics are discussed in [20].
@@
The dependence of the optimal success probability on $N$ is quite revealing. One might expect that increasing the length of the string makes edge detection harder, since the number of possible locations of the edge increases. However, we observe that for sufficiently large strings, the success probability is actually an increasing function of $N$. This behavior can be intuitively understood by noticing that larger domains potentially provide more information about the state of the particles they consist of, making edge detection easier. This tradeoff results in an increasing success probability. This is in contrast with the QCP, for which the success probability decreases with $N$. Simulations validate this intuition: We have compelling numerical evidence that a two-step strategy, involving the estimation of the known state of each domain from a vanishing fraction of particles (e.g., of the order of $\sqrt{N}$ ), followed by the application of the QCP protocol for edge localization, is asymptotically optimal.
@@ 3.1 Numerical results
The green open circles in Fig. 1 illustrate the results of our calculation, obtained using Eqs. (15) and (16) for $d=2$ (qubits). Similar trends are observed for other values of $d$. These points represent the ultimate limit to the achievable success probability in detecting the edge of a domain along a string of quantum particles.
@@
In simple terms, the joint probability of obtaining $\lambda$ and succeed amounts to the sum of the squares of the diagonal entries of $\sqrt{{\sf G}^{\lambda}}$. The success probability $P^{\rm SRM}_{\rm s}=\sum_{\lambda}P^{\rm SRM}_{\lambda}$ gives a lower bound to $P_{\rm s}$ since it results from a particular choice for ${\sf E}_{k}$ in Eq. (15).
Since no optimization is required to compute $P^{\rm SRM}_{\lambda}$, it becomes feasible to evaluate the success probability $P^{\rm SRM}_{\rm s}$ for significantly larger values of $N$. Additionally, the SRM equips us with the tools to derive analytical results for the asymptotic behavior of $P_{\rm s}$ as $N$ grows large, a topic we discuss in the next subsection.
In Fig. 1, the pink open circles, connected by lines, depict the results of our numerical evaluation of $P^{\rm SRM}_{\rm s}$ for $d=2$ (similar patterns hold for other values of $d$). It is noteworthy that the pink circles differ almost imperceptibly from the green ones, which represent the optimal protocol, indicating that the SRM is nearly optimal. Both sets of points fall underneath the dashed line but gradually approach it. This line represents a scenario where the quantum states of the domains are known and this information is used for edge detection. More precisely, it depicts the average success probability over $|\phi_{0}\rangle$ and $|\phi_{1}\rangle$ in Eq. (2) when the measurement is allowed to depend on these states and $N\to\infty$. This average, that we denote as $P_{\rm s}^{\rm known}$, can be easily computed (see below) using the results from the QCP derived in [16] and also serves as an upper bound to $P_{\rm s}$ in the scenario considered in this paper. The approaching of the pink circles to this line also showcases the asymptotic optimality of the SRM.
@@ 4 Known state in one domain
Here, $n_{0}+n_{1}=k$, with $0\leq n_{0}\leq k$. The set $\{P_{\sigma}\}$ stands for the unitary representation of $S_{k}$ on $({\mathbb{C}}^{2})^{k}$, which permutes the $k$ particles, and the sum extents only to those permutations $\sigma$ that do not leave the state $|0\rangle^{\otimes n_{0}}\otimes|1\rangle^{\otimes n_{1}}$ invariant. Clearly, only states with the same number $n_{1}$ may exhibit a non-vanishing overlap. This class of bases, known as Jordan bases [40], play a key role in programmable discriminators [41, 42].
@@ Appendix A Schur basis and Gram matrices
$\left|{\Phi_{\alpha}}\right\rangle=\sum_{\lambda}\left|{V_{\lambda}(\phi_{0},\phi_{1})}\right\rangle\otimes|\Omega_{\alpha}^{\lambda}\rangle\,.$ (34)
@@ Appendix B Asymptotic results
As explained in Section 3.2, we introduce a scaled version of $j$, defined as $x=j/(N/2)$, and compute the Maclaurin series expansion of the joint probability $P(x):=P^{\rm SRM}_{\lambda(x)}$ (indices are dropped to avoid clutter) up to a high order in $x$. At the leading order in $1/N$, the joint probability is given by the series:
@@
The (exact) coefficients $a_{r}$ have been computed up to $r=15$ or $r=18$, depending on $d$, using Wolfram Mathematica 13.3. For $d=2,3,4$, and $8$, they are collected in Tables 36.

### v2

@@ 1 Introduction
In a recent paper [19], the concept of unsupervised classification of quantum data was introduced, focusing on systems with unknown quantum states. The paper presents an optimal single-shot protocol for binary classification, capable of automatically sorting disordered input arrays with minimal assumptions while preserving partial information. It is demonstrated that this protocol surpasses classical methods for dimensionality three or higher. Furthermore, while the quantum data classification protocol runs efficiently on a quantum computer, the equivalent task with classical data involves solving an NP-complete problem. Its potential applications in high-energy physics are discussed in [20].
@@
The dependence of the optimal success probability on $N$ is quite revealing. One might expect that increasing the length of the string makes edge detection harder, since the number of possible locations of the edge increases. However, we observe that for sufficiently large strings, the success probability is actually an increasing function of $N$. This behavior can be intuitively understood by noticing that larger domains potentially provide more information about the state of the particles they consist of, making edge detection easier. This tradeoff results in an increasing success probability. This is in contrast with the QCP, for which the success probability decreases with $N$. Simulations validate this intuition: We have compelling numerical evidence that a two-step strategy, involving the estimation of the known state of each domain from a vanishing fraction of particles (e.g., of the order of $\sqrt{N}$ ), followed by the application of the optimal (non-local) QCP protocol for edge localization, is asymptotically optimal.
@@ 3.1 Numerical results
The green points in Fig. 1 illustrate the results of our calculation, obtained using Eqs. (15) and (16) for $d=2$ (qubits). Similar trends are observed for other values of $d$. These points represent the ultimate limit to the achievable success probability in detecting the edge of a domain along a string of quantum particles.
@@
In simple terms, the joint probability of obtaining $\lambda$ and succeed amounts to the sum of the squares of the diagonal entries of $\sqrt{{\sf G}^{\lambda}}$. The success probability $P^{\rm SRM}=\sum_{\lambda}P^{\rm SRM}_{\lambda}$ gives a lower bound to $P_{\rm s}$ since it results from a particular choice for ${\sf E}_{k}$ in Eq. (15).
Since no optimization is required to compute $P^{\rm SRM}_{\lambda}$, it becomes feasible to evaluate the success probability $P^{\rm SRM}$ for significantly larger values of $N$. Additionally, the SRM equips us with the tools to derive analytical results for the asymptotic behavior of $P_{\rm s}$ as $N$ grows large, a topic we discuss in the next subsection.
In Fig. 1, the pink points, connected by lines, depict the results of our numerical evaluation of $P^{\rm SRM}$ for $d=2$ (similar patterns hold for other values of $d$). It is noteworthy that the pink points differ almost imperceptibly from the green ones, which represent the optimal protocol, indicating that the SRM is nearly optimal. Both sets of points fall underneath the dashed line but gradually approach it. This line represents a scenario where the quantum states of the domains are known and this information is used for edge detection. More precisely, it depicts the average success probability over $|\phi_{0}\rangle$ and $|\phi_{1}\rangle$ in Eq. (2) when the measurement is allowed to depend on these states and $N\to\infty$. This average, that we denote as $P_{\rm s}^{\rm known}$, can be easily computed (see below) using the results from the QCP derived in [16] and also serves as an upper bound to $P_{\rm s}$ in the scenario considered in this paper. The approaching of the pink points to this line also showcases the asymptotic optimality of the SRM.
@@ 4 Known state in one domain
Here, $n_{0}+n_{1}=k$, with $0\leq n_{0}\leq k$. The set $\{P_{\sigma}\}$ stands for the unitary representation of $S_{k}$ on $({\mathbb{C}}^{2})^{k}$ that permutes the $k$ particles and the sum extents only to those permutations $\sigma$ that do not leave the state $|0\rangle^{\otimes n_{0}}\otimes|1\rangle^{\otimes n_{1}}$ invariant. Clearly, only states with the same number $n_{1}$ may exhibit a non-vanishing overlap. This class of bases, known as Jordan bases [40], play a key role in programmable discriminators [41, 42].
@@ 5 Summary and conclusions
Another important question that remains to be addressed in depth concerns the robustness of our protocol against imperfections in and along the string of particles. These may arise in various ways, e.g., the states may not be perfectly pure or may be slightly different along the string within each domain. A detailed analysis of this problem requires additional techniques beyond those introduced in this work and is currently under investigation. However, under reasonable assumptions [44, 45], the domains’ states may be thought of as if they undergo a small depolarization, in which case deviations from our results would be proportional to the depolarizing parameter, supporting the conclusion that our edge detection protocol, as it stands, is robust.
@@ Appendix A Schur basis and Gram matrices
$\left|{\Phi_{\alpha}}\right\rangle=\sum_{\lambda}\left|{V_{\lambda}(\phi_{0},\phi_{1})}\right\rangle\otimes|\Omega_{\alpha}^{\lambda}\rangle,$ (34)
@@ Appendix B Asymptotic results
As explained in Section 3.2, we introduce a scaled version of $j$, defined as $x=j/(N/2)$, and compute the Maclaurin series expansion of the joint probability $P(x):=P^{\rm SRM}_{\lambda(x)}$ (indices are dropped to avoid clutter) up to a high order, $2R$, in $x$. At the leading order in $1/N$, the joint probability is given by the series:
@@
The (exact) coefficients $a_{r}$ have been computed up to $r=15$ or $r=18$, depending on $d$, using Wolfram Mathematica 13.3, with the corresponding code available in [48]. For $d=2,3,4$, and $8$, they are collected in Tables 36.